# TASER: Table Agents for Schema‑guided Extraction and Recommendation

### v1

@@ 4 Experimental Setup
All experiments use gpt-4o-2024-11-20 as the LLM.
4.1 Evaluation Metrics
Detection Metrics: We report recall, precision, F1, and accuracy for table detection, prioritizing recall to eliminate the risk of missed financial holdings tables in extraction.
Extraction Metrics: We also assess extraction completeness by comparing TASER’s outputs to the ground truth labels. Each Financial Holdings Table houses a total net asset value, which we manual label for . We compare our labeled ground truth with the addition of all our extractions - this will be dubbed the total absolute difference (TAD).
Schema Refinement Metrics: Coverage is the fraction of unmatched holdings aligned with at least one schema suggestion, using RapidFuzz string similarity with a lenient ($\geq 70$) threshold. We also report the number of new matched holdings after re-extraction with the suggested schemas added to Portfolio. Diversity is the average pairwise Levenshtein distance between suggestion attributes (name and generated schema). Collision rate denotes the proportion of duplicate schemas among suggestions (reported on name and generated schema).
Dataset: We curate a diverse corpus of financial filings totaling 22,584 pages, 28M tokens, and \$731.7B in holdings. Among 3,213 tables, 57.53% exhibit hierarchical structure (via spanning cells). All Portfolio of Investment tables (100%) are hierarchical. While 39% of portfolios are single-page, 60.2% span multiple pages. The average length is 3.24 pages ($\sigma=3.41$, max = 19). This variability underscores the need for multi-page detection and consolidation.
@@ 5.1 Quantitative Evaluation
Detection: Table 2 shows that all TASER ablations achieve perfect recall ($\sim$100%), but precision ranges from 32.8% (Table Transformer) up to 43.4% (Full Schema Prompting), driving F1 scores between 49.3% and 59.4%. Embedding the full Portfolio schema in the prompt boosts precision by over 10% relative to the vision‑only baseline and yields the highest F1 (59.4%) and accuracy (66.4%), demonstrating that in‑context schema guidance is critical for filtering true holdings tables from noisy pages.
Extraction: Table 3 confirms that schema anchored extraction improves dollar‑value fidelity. Full Schema Prompting attains the lowest total absolute difference (\$102.8M) and the smallest unaccounted share (0.014%), outperforming Raw Text Prompting (\$107.1M, 0.015%) and Structured CoT (\$120.6M, 0.016%). Direct Schema Application (skipping detection) incurs a higher error (\$118.9M; 0.016%) by parsing spurious non‑holding pages. Therefore, full-schema detection followed by schema-driven extraction filters to true holdings tables, tightening dollar-value estimates.
@@ 5.2 Success Highlights
Cross-Document Consistency: TASER classifies and extracts holdings tables with varying titles (e.g., ”Portfolio of Investments”, ”Schedule of Holdings”, or ”Investment Portfolio”) and diverse structural formats. Despite the immense complexity of inputs, TASER consistently extracts and transforms the tables, ensuring that the final output appears as if sourced from a uniform set.
@@
Contextual Understanding: TASER excels in handling contextual nuances within financial documents, such as interpreting negative values denoted by parentheses (e.g., (140)) in zero-shot settings. Such domain-specific attributes are important in understanding financial tables.
@@
Extracting Intricate Semantics: TASER demonstrates a strong semantic understanding of financial terminology, which empowers it to not only extract but also correctly comprehend the financial information it processes. For instance, TASER adeptly parsed the table entry “GBP 4,700,000 — UK Treasury 0% 19/02/2024 — 4668 — 1.48,” correctly identifying the holding as a bond and extracting its attributes: quantity, market value, coupon rate, maturity date, and issuer.
@@
Figure 6: Reduction in total absolute difference (TAD) after resolving unmatched holdings. Remaining TAD is calculated after sequential reconciliation of unmatched holdings. Most error reduction is achieved by resolving the most significant unmatched holdings; additional reconciliation yields only incremental improvement, consistent with a heavy-tailed value distribution (Figs. 8 & 9).
5.3 Batch Size Tradeoffs in Schema Refinement
@@
Figure 5 (left) reveals that larger batch sizes (250, 500) rapidly expand the schema set in the first few iterations, achieving faster initial coverage as measured by iterations. However, this early acceleration comes at the cost of quickly reaching saturation, after which few new unique schemas are discovered. In contrast, smaller batch sizes (10, 50, 100) require more iterations to reach the same number of unmatched holdings seen, but continue yielding new unique schemas for much longer, resulting in the highest overall diversity when normalized by data processed (Figure 5, right). This improvement in coverage for smaller batches is offset by increased redundancy. As shown in Appendix 7, smaller batches incur substantially more duplicate or overlapping schema suggestions, reflecting a more granular and exploratory nature of their schema search. Overall, these results highlight a key tradeoff in agentic schema refinement: larger batches accelerate early-stage discovery but plateau quickly, while smaller batches maximize cumulative schema diversity at the cost of higher redundancy and computation. Our results establish that schema refinement via agentic suggestion and feedback is both tractable and tunable: by adjusting batch size, practitioners can directly control the balance between coverage and efficiency, guiding schema evolution to maximize actionable suggestions while minimizing redundancy and annotation waste. This provides a principled, quantitative framework for automated schema refinement in noisy, open-ended domains
@@
Schema Diversity and Utilization: Schema diversity, as measured by the average pairwise Levenshtein distance, is maximized for moderate batch sizes (100–250), as shown in Appendix K. While larger batch sizes (500) yield a higher proportion of utilized schemas—up to 59%—smaller, more diverse batches tend to have lower utilization rates (Table 9). Furthermore, the accretive gain in 402 additional unique schemas yielded only marginal improvements in holding coverage (6.1%). Figure 10 illustrates this tradeoff: smaller batch sizes cover more unmatched holdings at the expense of efficiency (96.1% coverage for 29.0% utilization at batch size 10), whereas larger batches achieve higher schema utilization (59.0% at batch size 500), a greater proportion of actionable suggestions. This spectrum allows TASER to target either exhaustive coverage or more efficient, high-value schema evolution, depending on application needs.
5.4 Schema Refinement Optimality
Systematic ablation of batch size during schema refinement for unmatched financial holdings demonstrates a tunable tradeoff between coverage—the absolute number of holdings addressed—and utilization—the efficiency of actionable schema suggestion. By selecting appropriate batch sizes, schema refinement can be tailored to maximize coverage, efficiency, or a balanced combination, supporting principled, automated schema evolution in real-world information extraction. Practically, these competing objectives can be summarized as follows: small batches maximize recall, ensuring that most unmatched entities are eventually addressed, while large batches drive high-precision, low-waste schema discovery. Practically, this means that a mixed or adaptive batching strategy may be optimal—using larger batches to quickly identify high-yield schemas, followed by smaller, targeted batches to exhaustively cover residual unmatched cases.
@@ 5.4 Schema Refinement Optimality
Improvements in TAD: Resolving the largest unmatched holdings yields a reduction in total absolute difference (TAD) of approximately 7–10% across batch sizes, with the majority of improvement achieved by reconciling just the top 10–20% of holdings (Table 5). Beyond this, further reconciliation produces only marginal gains, reflecting the highly heavy-tailed nature of recoverable value.
@@ 6 Conclusion
By combining LLM agents with dynamic, schema‑anchored prompting and recursive validation loops, TASER offers a new proposal for extracting and normalizing complex holdings tables from raw financial filings. TASER’s high precision and recall across diverse financial layouts underscore the potential of agentic, domain‑aware reasoning for scalable and accurate document understanding.
Appendix A Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates (”JPMorgan”) and is not a product of the Research Department of JPMorgan. JPMorgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Appendix B Limitations
@@
Appendix C Camelot Table Parsing Modes
@@
Appendix D TASER Annotation Process
@@
Appendix E TASER Dataset Release
@@
Appendix F Document Preprocessing
@@
Appendix G Parallelization and Fund Construction
@@
Dataset Avg. # Tables Avg. Rows Avg. Columns Avg. Spanning
Per Topology Per Table Per Table Cells per Table
@@
Dataset # Unique Cell Avg. # Tables Avg. Rows Avg. Columns Maximum Page Span
# Tables Topologies Per Topology Per Table Per Table
@@
Appendix H Schema Definitions and Portfolio Model
@@
Appendix I Ablation Strategies
@@
Appendix J Aggregation and Conflict Resolution of Schema Suggestions
@@
Appendix K Schema Suggestion Diversity
@@
Appendix L Example Holdings Tables
@@
TASER
(a) Raw Text Prompting 1,495 0.33 5,414 20.37 6,909 20.69
@@
Figure 12: Detection prompts for all ablation strategies. Each section is labeled with its corresponding ablation strategy
Instrument Category Count Example
Equities 28,737 Taiwan Semiconductor Manufacturing
Debt 17,105 US Treasury 4.69% 09/05/2024
Unmatched Instruments 16,822 EUR
Forwards 8,023 Bought USD Sold KRW at 0.00072513
Options 977 Written Call Unilever 4050
Futures 720 US 5 Year Bond Future
Swaps 776 Pay fixed 3.026% receive float. (1d SOFR)
ELNs 292 BNP (Laobaixing Pharm. Chain (A)) ELN 22/07/2024
@@
8 isin: Optional[str] = Field(None, description="International Securities Identification Number")
@@
10 description: Optional[str] = Field(None, description="Description or name of the instrument")
@@
13
14class Equity(BaseInstrument):
@@
17
18class Option(BaseInstrument):
@@
22 expiration_date: Optional[datetime] = Field(None, description="Expiration date of the option")
@@
24
25class Swap(BaseInstrument):
@@
28 fixed_rate: Optional[float] = Field(None, description="Fixed rate component (if applicable)")
29 floating_rate_index: Optional[str] = Field(None, description="Index used for floating rate leg")
@@
32
33class Forward(BaseInstrument):
@@
36 settlement_date: Optional[datetime] = Field(None, description="Settlement date for the forward")
37
38class Future(BaseInstrument):
@@
41 expiration_date: Optional[datetime] = Field(None, description="Expiration date of the future")
42
43class Debt(BaseInstrument):
@@
45 coupon_rate: Optional[float] = Field(None, description="Annual coupon rate of the debt/bond")
46 maturity_date: Optional[datetime] = Field(None, description="Maturity date of the debt/bond")
@@
48
49class EquityLinkedNote(BaseInstrument):
@@
2 description: str = Field(..., description="Text of the unknown instrument.")
3 name: str = Field(..., description="Suggested classification of the description or type")
4 market_value: Optional[float] = Field(None, description="Market value associated with the instrument")
5
6class Portfolio(BaseModel):
7 fund_name: Optional[str] = Field(None, description="Name of the fund that the portfolio belongs to")
8 value_in_thousands: bool = Field(False, description="True if the market value is based on thousands")
9 equities: Optional[List[Equity]] = Field(default_factory=list, description="List of equities")
10 options: Optional[List[Option]] = Field(default_factory=list, description="List of options")
11 swaps: Optional[List[Swap]] = Field(default_factory=list, description="List of swaps")
12 forwards: Optional[List[Forward]] = Field(default_factory=list, description="List of forwards")
13 futures: Optional[List[Future]] = Field(default_factory=list, description="List of futures")
14 debt: Optional[List[Debt]] = Field(default_factory=list, description="List of debt instruments")
15 elns: Optional[List[EquityLinkedNote]] = Field(default_factory=list, description="List of equity linked notes")
16 other_instruments: Optional[List[Other]] = Field(default_factory=list, description="The list of instruments that do not match any other type")
@@
13- Given the current schema (JSON below), propose new classes or modifications so each holding can be classified.
14- If a holding matches a previously suggested class, propose new optional fields if needed.
15- Return your schema suggestions as a list of Pydantic SchemaSuggestion model objects.
16
17Current Portfolio Schema:
18{Portfolio.model_json_schema()}
19
20Batch of unmatched holdings:
21{unmatched_holdings[start : start + batch_size]}
22
23Previously seen suggestions (optional, from prior batches):
24{previous_suggestions if previous_suggestions else None}
25
26For each unique holding, propose:
27- A new schema class, or a modification to an existing class (add or refine fields).
28- Specify all required and optional fields with Python type hints.
29- If similar to an earlier suggestion, mark only new fields as optional.
30- Provide a sample match (the original holding string).
31- Output format: a Python list of SchemaSuggestion objects, as defined below.
32"""
33
@@
35class SchemaSuggestion(BaseModel):

### v4

@@ 4 Experimental Setup
Detection Metrics: We report recall, precision, F1, and accuracy for table detection, prioritizing recall to avoid missing Financial Holdings Tables.
Extraction Metrics: We assess extraction completeness by comparing TASER’s outputs to ground truth labels. We manually label a total net asset value for each Holdings Table. We then compare this ground truth with our extractions, dubbed the total absolute difference (TAD).
Schema Refinement Metrics: Coverage is the fraction of unmatched holdings aligned with at least one schema suggestion, using RapidFuzz string similarity with a lenient ($\geq 70$) threshold. We also report the number of new matched holdings after re-extraction with the suggested schemas added to Portfolio. Diversity is the average pairwise Levenshtein distance between suggestion attributes (name and generated schema). Collision rate denotes the proportion of duplicate suggestions.
Dataset and Model: We curate a diverse corpus totaling 22,584 pages, 28M tokens, and \$731.7B in holdings. Among 3,213 tables, 57.53% exhibit hierarchical structure (via spanning cells). All Holdings Tables (100%) are hierarchical. While 39% of portfolios are single-page, 60.2% span multiple pages. The average length is 3.24 pages ($\sigma=3.41$, max = 19). This variability underscores the need for multi-page detection and consolidation. Unless explicitly stated otherwise, all experiments use gpt-4o-2024-11-20 as the LLM.
@@ 5.1 Quantitative Evaluation
Detection: Table 2 shows that all TASER ablations achieve perfect recall ($\sim$100%), but precision ranges from 32.8% (Table Transformer) up to 43.4% (Full Schema Prompting), driving F1 scores between 49.3% and 59.4%. Embedding the full Portfolio schema in the prompt boosts precision by over 10% relative to the vision‑only baseline and yields the highest F1 (59.4%) and accuracy (66.4%), demonstrating that in‑context schema guidance is critical.
Extraction: Table 2 confirms schema-anchored extraction improves dollar‑value fidelity. Full Schema Prompting attains the lowest absolute difference (\$102.8M) and smallest unaccounted share (0.014%), outperforming Raw Text Prompting (\$107.1M, 0.015%) and Structured CoT (\$120.6M, 0.016%). Direct Schema Application (skipping detection) incurs a higher error (\$118.9M; 0.016%) by parsing spurious non‑holding pages.
@@ 5.2 Success Highlights
Cross-Document Consistency: TASER classifies and extracts Holdings Tables despite varying titles (e.g., "Portfolio of Investments", "Schedule of Holdings", or "Investment Portfolio") and diverse structural formats. Despite the immense complexity of inputs, TASER consistently extracts and transforms these tables, ensuring that the final output appears as if sourced from a uniform set.
@@
Contextual Understanding: TASER excels in handling contextual nuances, such as interpreting negative values denoted by parentheses (e.g., (140)) in zero-shot settings. Such domain-specific attributes are important for financial tables.
@@
Extracting Intricate Semantics: TASER demonstrates a strong semantic understanding of financial terminology, which empowers it to extract accurately. For instance, TASER adeptly parsed the table entry “GBP 4,700,000 | UK Treasury 0% 19/02/2024 | 4668 | 1.48,” correctly identifying the holding as a bond and extracting its attributes: quantity, market value, coupon rate, maturity date, and issuer.
@@
Figure 6: Reduction in TAD after resolving unmatched holdings. Remaining TAD is calculated after sequential reconciliation of unmatched holdings. Error reduction is achieved by resolving the most significant unmatched holdings (Figs. 8 & 9).
5.3 Batch Size Tradeoffs in Refinement
@@
Figure 5 (left) reveals that larger batch sizes (250, 500) rapidly expand the schema. However, this early acceleration comes at the cost of early saturation, after which few new unique schemas are discovered. In contrast, smaller batches require more iterations to reach the same number of unmatched holdings seen, but continue yielding unique schemas, resulting in the highest diversity when normalized by data processed (Figure 5, right). This improvement in coverage, however, is offset by increased redundancy. As shown in Appendix Figure 7, smaller batches incur substantially more overlapping suggestions, reflecting a more granular and exploratory nature. Overall, these results highlight a key tradeoff: larger batches accelerate early discovery but plateau quickly, while smaller batches maximize cumulative schema diversity at the cost of redundancy and computation. Our results establish that schema refinement via agentic feedback is both tractable and tunable. This indicates that an adaptive batching strategy may be optimal: using larger batches to quickly identify high-yield schemas, followed by smaller batches for exhaustive diversity.
@@
Schema Diversity and Utilization: Schema diversity, as measured by the average pairwise Levenshtein distance, is maximized for moderate batch sizes (100–250), as shown in Appendix J. While larger batch sizes (500) yield a higher proportion of utilized schemas—up to 59%—smaller, more diverse batches tend to have lower utilization rates (Table 9). Furthermore, the accretive gain in 402 additional unique schemas yielded only marginal improvements in holding coverage (6.1%). Figure 10 illustrates this tradeoff: smaller batch sizes cover more unmatched holdings at the expense of efficiency (96.1% coverage for 29.0% utilization at batch size 10), whereas larger batches achieve higher schema utilization (59.0% at batch size 500).
@@
Improvements in TAD: Resolving the largest unmatched holdings yields a reduction in TAD of approximately 7–10% across batch sizes, with the majority of improvement achieved by reconciling just the top 10–20% of holdings (Table 5).
5.4 Deployment of TASER
We outline the deployment architecture of TASER in Figure 3, with additional system architecture details provided in Appendix A.
@@ 6 Conclusion
We present TASER for extracting complex Holdings Tables from documents through continual learning. Our high precision and recall across diverse layouts underscore the potential of agentic continual learning for financial table extraction.
Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates ("JPMorgan”) and is not a product of the Research Department of JPMorgan. JPMorgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Limitations
@@
Appendix A TASER System Architecture
A.1 Overview
TASER is an event-driven microservice architecture composed of several semi-independent agents, such as text extraction and table detection modules. Each agent operates on a shared cloud infrastructure, which typically consists of an AWS Lambda function that listens to a queue (Amazon SQS), manages file operations in Amazon S3, and utilizes a dead-letter queue (DLQ) for retry handling.
A.2 User Interaction and Deployment
Users may interact with TASER either through a graphical user interface (UI) or directly via the application programming interface (API). Both the UI and API are deployed within an Amazon ECS cluster, with horizontal and vertical scaling enabled to respond to user demand. The UI is updated by periodically polling the API for task status, which is maintained in Amazon DynamoDB.
A.3 Task Cache
An additional component is the task cache, implemented using DynamoDB. The primary purpose of the cache is to reduce latency and operational costs by retrieving previously computed results. For all agents, the cached value is typically an S3 location indicating where the result of a prior execution is stored. The cache key varies by agent but generally includes model-specific information (e.g., model name, temperature settings) and the prompt used for generation. A cache entry for the table detection agent is structured as follows:
Key: {schema_id, prompt_id, document_text_id,
model_name, model_configs}
Value: S3 location
A.4 Design Rationale: API-Centric Orchestration
A core design decision for TASER was to utilize an API-centric approach for orchestrating workflows. An alternative would have been to configure each agent to automatically trigger an AWS Lambda function upon the arrival of a new file in S3 (either by directly configuring the S3 bucket or using AWS EventBridge). While this method is feasible, it means that any change to workflows would require reprovisioning AWS infrastructure. In contrast, with the current architecture, redesigning a workflow is as simple as refactoring the UI or API client code (for example, changing the order of API calls) and redeploying the application, without modifying the underlying cloud infrastructure.
Appendix B Camelot Table Parsing Modes
@@
Appendix C TASER Annotation Process
@@
Appendix D TASER Dataset Release
@@
Appendix E Document Preprocessing
@@
Appendix F Parallelization and Fund Construction
@@
Avg. # Tables Avg. Rows Avg. Columns Avg. Spanning
Dataset Per Topology Per Table Per Table Cells per Table
@@
# Unique Cell Avg. # Tables Avg. Rows Avg. Columns Maximum Page Span
Dataset # Tables Topologies Per Topology Per Table Per Table
@@
Appendix G Schema Definitions and Portfolio Model
@@
Appendix H Ablation Strategies
@@
Appendix I Aggregation and Conflict Resolution of Schema Suggestions
@@
Appendix J Schema Suggestion Diversity
@@
Appendix K Example Holdings Tables
@@
TASER (a) Raw Text Prompting 1,495 0.33 5,414 20.37 6,909 20.69
@@
Instrument Category Count Example
Equities 28,737 Taiwan Semiconductor Manufacturing
Debt 17,105 US Treasury 4.69% 09/05/2024
Unmatched Instruments 16,822 EUR
Forwards 8,023 Bought USD Sold KRW at 0.00072513
Options 977 Written Call Unilever 4050
Futures 720 US 5 Year Bond Future
Swaps 776 Pay fixed 3.026% receive float. (1d SOFR)
ELNs 292 BNP (Laobaixing Pharm. Chain (A)) ELN 22/07/2024
@@
Figure 12: Detection prompts for all ablation strategies. Each section is labeled with its corresponding ablation strategy.
@@
8 isin: Optional[str] = Field(None,
9 description="International Securities Identification Number")
@@
11 description: Optional[str] = Field(None,
12 description="Description or name of the instrument")
@@
15
16class Equity(BaseInstrument):
@@
19
20class Option(BaseInstrument):
@@
24 expiration_date: Optional[datetime] = Field(None,
25 description="Expiration date of the option")
@@
27
28class Swap(BaseInstrument):
@@
31 fixed_rate: Optional[float] = Field(None,
32 description="Fixed rate component (if applicable)")
33 floating_rate_index: Optional[str] = Field(None,
34 description="Index used for floating rate leg")
@@
37
38class Forward(BaseInstrument):
@@
41 settlement_date: Optional[datetime] = Field(None,
42 description="Settlement date for the forward")
43
44class Future(BaseInstrument):
@@
47 expiration_date: Optional[datetime] = Field(None,
48 description="Expiration date of the future")
49
50class Debt(BaseInstrument):
@@
52 coupon_rate: Optional[float] = Field(None,
53 description="Annual coupon rate of the debt/bond")
54 maturity_date: Optional[datetime] = Field(None,
55 description="Maturity date of the debt/bond")
@@
57
58class EquityLinkedNote(BaseInstrument):
@@
2 description: str = Field(...,
3 description="Text of the unknown instrument.")
4 name: str = Field(...,
5 description="Suggested classification of the description or type")
6 market_value: Optional[float] = Field(None,
7 description="Market value associated with the instrument"
8 )
9
10class Portfolio(BaseModel):
11 fund_name: Optional[str] = Field(None,
12 description="Name of the fund that the portfolio belongs to")
13 value_in_thousands: bool = Field(False,
14 description="True if the market value is based on thousands")
15 equities: Optional[List[Equity]] = Field(default_factory=list,
16 description="List of equities")
17 options: Optional[List[Option]] = Field(default_factory=list,
18 description="List of options")
19 swaps: Optional[List[Swap]] = Field(default_factory=list,
20 description="List of swaps")
21 forwards: Optional[List[Forward]] = Field(default_factory=list,
22 description="List of forwards")
23 futures: Optional[List[Future]] = Field(default_factory=list,
24 description="List of futures")
25 debt: Optional[List[Debt]] = Field(default_factory=list,
26 description="List of debt instruments")
27 elns: Optional[List[EquityLinkedNote]] = Field(default_factory=list,
28 description="List of equity linked notes")
29 other_instruments: Optional[List[Other]] = Field(default_factory=list,
30 description="The list of instruments that do not match any other type")
@@
13- Given the current schema (JSON below), propose new classes or modifications so each holding
14 can be classified.
15- If a holding matches a previously suggested class, propose new optional fields if needed.
16- Return your schema suggestions as a list of Pydantic SchemaSuggestion model objects.
17
18Current Portfolio Schema:
19{Portfolio.model_json_schema()}
20
21Batch of unmatched holdings:
22{unmatched_holdings[start : start + batch_size]}
23
24Previously seen suggestions (optional, from prior batches):
25{previous_suggestions if previous_suggestions else None}
26
27For each unique holding, propose:
28- A new schema class, or a modification to an existing class (add or refine fields).
29- Specify all required and optional fields with Python type hints.
30- If similar to an earlier suggestion, mark only new fields as optional.
31- Provide a sample match (the original holding string).
32- Output format: a Python list of SchemaSuggestion objects, as defined below.
33"""
@@
35
36class SchemaSuggestion(BaseModel):